OpenStudy (anonymous):

What is the phase shift for the following? a. A sine wave with the maximum amplitude at time zero b. A sine wave with maximum amplitude after 1/4 cycle c. A sine wave with zero amplitude after 3/4 cycle and increasing

OpenStudy (anonymous):

a)\(\phi=-\pi/2\) b) \(\phi=-\pi/4\) c) \(\phi=-\pi/4\)

OpenStudy (anonymous):